Only need solve the underline parts.please give full explanations.Why and how many terms might required.  Other parts are DONE. 1. Equation 12 shows two expressions for a magnetic field produced by a current carrying wire. One of them is approximated by
Answer #1

Given that

Assume L>>x(Bapprox(x))

L>>(Bexact(x))

x=L/10 to 2L

L=I=1

  311_(19.1) 03 on = 311(10:0 +1 ) og det en an (bo nie + ) 03 Home wife (@) ***) F e = (2) fora [1=7:.] = 9-X () Xopydog (03-.2.0(0.9805806) Besact(*) = x 19.611 --a) Bappsor (*) = Bapprox (N) = x 20 - (b) i At x=qL, x=2 (²2=1) (1+4x2?)\2 Bexact (N) =So, at x=1 Ho 2% (+4) - - 4T L5 Mo 2I , Bexact(x) = ? Bappogox = | behact - bеff%ss } 1 dislebence - X (Bexact + Bappxox)/2 1Bapprox (2) x Mo fro HTT Ba 2/2 1.8 2.0 N Bexact x Mozof Ba 2 Xitha? 0.2 0. 6 0 1.4 1.9 2.0 Mot a nd good approximation fox x2! ги (у ця)/2 Bionomias expansion of (14aj12 = 1 (2) + (-) (-) 4*, (ət:-) ( ) ... (1+ay - - - - - - - . ( ) - 2 ( ) - 2 Ця? + 6x4 at xav, BE No (1+1+6) Berminyak (6) which is not Valve oblined same 98 as va eme Catliex. - which was B so, for teamsBappbox (x) - Mo 201 HT Bionomial expansion expression, B(x) = more I ( 1+ ux%, 2)\/2 at x= 2/10, Bla) = no 2.176 (174(+)?/12